OpenStudy (anonymous):

PLEASE HELP I WILL GIVE FAN AND MEDAL Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 12 + (6 ÷ 3) = 18 ÷ 3 = 6 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. 32 + 8 ÷ 4 = 40 ÷ 4 = 10 D. (20 – 5) • 3 = 15 • 3 = 45 Please help me @mathman101

OpenStudy (anonymous):

What do you think? Hint: you divide and multiply before you add or subtract.

OpenStudy (anonymous):

Maybe D and A?

OpenStudy (anonymous):



OpenStudy (anonymous):

@troller not funny

OpenStudy (anonymous):

In A you first divide: 6/3 = 2, so you get: \[12+2=18/3=6\] this is not true. But you are right about D, try look at B and C again, one of them is also correct.

OpenStudy (anonymous):

I think its C, maybe?

OpenStudy (anonymous):

You divide before adding, so in C you first divide 8 ÷ 4 = 2, and then you get \[32+2=40÷4\] is that correct?

OpenStudy (anonymous):

B. and D.

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Order_of_operations

OpenStudy (anonymous):

Thanks @Tommynaiter